OpenStudy (meilendurcer):

Find the length of the hypotenuse of a right triangle with legs of 9 cm and 12 cm.

OpenStudy (anonymous):

for this problem youre going to want to use a^2 + b^2 = c^2 (Pythagorean theorem)

OpenStudy (anonymous):

a^2+b^2=c^2 9^2+12^2=c^2

OpenStudy (anonymous):

81+144=c^2

OpenStudy (anonymous):

c^2=228

jhonyy9 (jhonyy9):

so this not is the end

jhonyy9 (jhonyy9):

what is c = ?

OpenStudy (anonymous):

she ll take the square root ofcourse

OpenStudy (anonymous):

\[c=\sqrt{228}\]

OpenStudy (anonymous):

c=15.1 cm